### v2

@@ A.5 Supplementary Analysis: Economic Analysis of PEM Electrolyzer Implementation
Response time ¡1 second Enables grid balancing revenue [s12]
@@
Optimized operation ¡\$1.00/kg Advanced materials, scale Industry target
@@ A.6.3 Uncertainty Quantification Details
A.6.4 Complete Physics Constraints
The physics-informed loss function incorporates eleven constraints total, with five auxiliary components complementing the six fundamental laws in the main text:
Water Activity and Content Dynamics
$\mathcal{L}_{\text{water\_activity}}=\left\|a-\frac{P_{H_{2}O}}{P_{\text{sat}}}\right\|^{2}$ (10)
$\mathcal{L}_{\text{water\_content}}=\left\|\lambda_{m}-\left(0.043+17.81a-39.85a^{2}+36.0a^{3}\right)\right\|^{2}$ (11)
where water activity $a$ determines membrane hydration state following Springer’s correlation.
Transport Property Constraints
$\mathcal{L}_{\text{diffusion}}=\left\|D_{\text{eff}}^{H_{2}}-\frac{\epsilon_{m}^{w}}{\tau_{m}}D_{H_{2},w}\right\|^{2}$ (12)
@@ A.6.4 Complete Physics Constraints
$\mathcal{L}_{\text{temperature\_dep}}=\left\|D_{H_{2},w}-D_{0}(1+\alpha_{T}(T-T_{\text{ref}}))\right\|^{2}$ (13)
Note: A linearized temperature dependence is employed instead of the full Arrhenius equation to enhance numerical stability during training while maintaining accuracy across the operating temperature range (25–85°C). The linear coefficient $\alpha_{T}=0.01$ K-1 was validated against experimental data.
Crossover Concentration
$\mathcal{L}_{\text{crossover}}=\left\|X_{H_{2}}^{\text{ca}}-\frac{\dot{M}_{H_{2}}^{\text{co}}}{\dot{M}_{O_{2}}}\times 100\%\right\|^{2}$ (14)
Implementation Details
The eleven physics constraints are implemented with equal weighting ($w_{j}=1$) in the composite loss function. The linearized temperature model (13) reduces computational complexity from $\mathcal{O}(n^{2})$ to $\mathcal{O}(n)$ compared to the exponential Arrhenius form while maintaining $R^{2}>0.99$ correlation with experimental diffusivity data in the operational range. This simplification enables stable gradient propagation through the 8-128-128-1 neural architecture without sacrificing prediction accuracy.
@@
  • Membrane porosity: $\epsilon_{m}^{w}=0.28$
  • Tortuosity factor: $\tau_{m}=1.5$
  • Darcy coefficient: $K_{d}=1.0\times 10^{-5}$ bar2/(A cm-2)
  • Reference temperature: $T_{\text{ref}}=298$ K
  • Base diffusion coefficient: $D_{0}=1.0\times 10^{-9}$ m2/s
These parameters were fixed based on literature consensus values to reduce model complexity while maintaining physical interpretability.
@@ A.7.2 Model Performance Reporting Convention
  • Rationale: Mean $\pm$ std provides robust performance estimate; best model demonstrates achievable upper bound

### v3

@@ A.5 Supplementary Analysis: Economic Analysis of PEM Electrolyzer Implementation
Response time <<1 second Enables grid balancing revenue [s12]
@@
Optimized operation <<\$1.00/kg Advanced materials, scale Industry target
@@ A.6.3 Uncertainty Quantification Details
A.6.4 Constitutive Physics Equations
Instead of imposing separate loss terms for each intermediate variable, the physics-informed framework calculates the target hydrogen crossover concentration $\Phi_{H_{2}}^{\text{phys}}$ by solving a system of coupled constitutive equations. These equations account for the material properties and transport phenomena as follows:
@@ A.6.4 Constitutive Physics Equations
The self-diffusion coefficient of hydrogen in water ($D_{H_{2},w}$) follows an Arrhenius relationship, with parameters fitted for the operating temperature range:
$D_{H_{2},w}(T)=7.734\times 10^{-6}\cdot\exp\left(-\frac{2225.4}{T}\right)\quad[\text{m}^{2}/\text{s}]$ (9)
where $T$ is the temperature in Kelvin.
Membrane Structural Parameters
Unlike simplified models that assume fixed structural parameters, our model dynamically calculates the water volume fraction ($\epsilon_{m,w}$) and tortuosity ($\tau_{m}$) based on the membrane’s specific water content ($\lambda_{m}$) and equivalent weight (EW):
$\epsilon_{m,w}=\frac{\lambda_{m}V_{w}}{V_{m}+\lambda_{m}V_{w}}$ (10)
$\tau_{m}=\left(1-0.85(1-\epsilon_{m,w})\right)^{-1}$ (11)
Here, $V_{w}$ is the molar volume of water and $V_{m}$ is the molar volume of the dry membrane. This formulation accurately captures the swelling effect of the membrane on effective diffusivity ($D_{\text{eff}}^{H_{2}}$).
Compression Effects on LGDL
The model incorporates the impact of cell assembly pressure on the Liquid/Gas Diffusion Layer (LGDL). The compressed thickness ($t_{\text{LGDL}}$) and permeability ($K_{\text{LGDL}}$) are corrected using a cubic relation with porosity:
$\phi_{\text{comp}}=1-\frac{t_{\text{initial}}}{t_{\text{comp}}}(1-\phi_{\text{initial}})$ (12)
$K_{\text{LGDL}}=K_{\text{initial}}\cdot\left(\frac{\phi_{\text{comp}}}{\phi_{\text{initial}}}\right)^{3}$ (13)
@@
The key parameters used in the physics model are summarized below. Note that empirical parameters $\alpha$ and $\beta$ are membrane-specific values optimized from experimental data.
  • Constants:
@@
    • Molar volume of water: $V_{w}=18.3$ cm3 mol-1
    • Hydrogen molar mass: $M_{H_{2}}=2.016\times 10^{-3}$ kg mol-1
  • Layer Properties (Default):
[…]
    • Ionomer volume fraction: $\epsilon_{io}=0.15$
    • LGDL Permeability: $K_{LGDL}=2.0\times 10^{-12}$ m2
    • CL Permeability: $K_{CL}=2.6\times 10^{-15}$ m2
  • Model Parameters (Example for Nafion 117):
    • Convection parameter $\alpha$: $5.52\times 10^{-3}$
    • Convection exponent $\beta$: Calculated via log-linear model ($a=0.440,b=0.0439$)
@@ A.7.2 Model Performance Reporting Convention
  • Rationale: Mean $\pm$ s.d. provides robust performance estimate; best model demonstrates achievable upper bound